The Reynolds Company buys from its suppliers on terms of 2/10, net 58. Reynolds has not been utilizing the discount offered and has been taking 70 days to pay its bills. The suppliers seem to accept this payment pattern, and Reynold’s credit rating has not been hurt.

Mr. Duke, Reynolds Company’s vice-president, has suggested that the company begin to take the discount offered. Mr. Duke proposes the company borrow from its bank at a stated rate of 11 percent. The bank requires a 11 percent compensating balance on these loans. Current account balances would not be available to meet any of this required compensating balance.

b. Calculate the annual rate of interest if the company borrows from the bank. (Use 365 days in a year. Do not round intermediate calculations. Round the final answer to 2 decimal places.)

Annual rate of Interest             %

Your bank will lend you $9,600 for 30 days at a cost of $102 interest.

a. What is your annual rate of interest? (Use 365 days in a year. Do not round intermediate calculations. Round the final answer to 2 decimal places.)

Annual rate of interest             %

b. What is your effective annual rate? (Use 365 days in a year. Do not round intermediate calculations. Round the final answer to 2 decimal places.)

Effective annual rate             %

Little Kimi Clothiers can borrow from its bank at 12 percent to take a cash discount. The terms of the cash discount are 2/18 net 85.

a. Compute the cost of not taking the cash discount. (Use 365 days in a year. Do not round the intermediate calculations. Round the final answer to 2 decimal places.)

Cost of not taking a cash discount             Not attempted%

The treasurer of Brandon Blue Sox is seeking a $30,000 loan for 180 days from the Brandon Credit Union. The stated interest rate is 15 percent and there is a 20 percent compensating balance requirement. The treasurer always keeps a minimum of $2,500 in the firm’s chequing account. These funds could count toward meeting any compensating balance requirements.

What is the annual rate of interest on this loan? (Use 365 days in a year. Do not round intermediate calculations. Round the final answer to 2 decimal places.)

Annual rate of interest             %
Answer #1

b) If the company borrows from the bank with 11% compensating balance requirement and 11% stated interest rate

On borrowing $100, 11% i.e.$11 will not be available for use and only $89 can be used. and an annual interest paid = 11% on $100 = $11

So, Annual Interest rate = $11/$89 = 12.36%

If the bank lends $9,600 for 30 days at a cost of $102 interest.

a) Rate of Interest for 30 days= 102/9600 =0.010625

Annual rate of interest = 0.010625* 365/30 = 0.129271 =12.93%

b. Effective annual rate = (1+30days interest rate)^(365/30) -1

=(1+0.010625)^(365/30) -1

=  0.137223

= 13.72%